# Learn Beyond The Answer: Training Language Models with Reflection for Mathematical Reasoning

### v1

@@ Abstract
Zhihan Zhang1thanks:   This work was done when Zhihan, Zhenwen, and Mengzhao were interns at Tencent AI Lab, Seattle., Zhenwen Liang1†, Wenhao Yu2, Dian Yu2, Mengzhao Jia1†, Dong Yu2, Meng Jiang1 1University of Notre Dame    2Tencent AI Lab, Seattle zzhang23@nd.edu
@@ 3 Approach
Data Annotation  Following a common approach Li et al. (2023); Yu et al. (2024); Li et al. (2024a), we employ an expert LM, GPT-4-turbo, to annotate the reflective sections for high-quality reasoning paths and minimal human effort [^2] We also tried LLaMA-3-70B for data annotation in Appendix A.1 but its performance lags behind GPT-4-turbo.. This entails reviewing the original problem and solution to generate a section consisting of the aforementioned two types of reflective reasoning. We prompt the expert model to choose between abstraction and analogy in follow-up reasoning based on the problem context. Figure 2 shows an annotated example with alternative reasoning and follow-up analogy, and the full annotation prompt is in Appendix E.
@@ 4.2.1 Tasks
Table 3: MathChat results compared with other open-source 7B math models. Baseline scores are from Liang et al. (2024). The best scores are bolded and the second bests are underlined. *Including both supervised fine-tuning and reinforcement learning data.
@@ 4.4.5 Stability of RefAug Data Annotation
In addition to the above perspectives, further analyses on the risk of data contamination and efficiency statistics in training and inference are presented in Appendix A.5 and A.6, respectively.
@@ 5 Conclusion
Some previous data augmentation studies in math reasoning created millions of data instances with OpenAI’s GPT models Li et al. (2024a); Tang et al. (2024); Huang et al. (2024b). While testing our method at a similar scale would be valuable, budget constraints limit our ability to do so. For instance, our augmentation data for MetaMath is capped at 40K instances. In Appendix A.1, we note that LLaMA-3-70B shows some promising performance in annotating RefAug data for math reasoning tasks, though its capabilities have not fully matched those of GPT-4 yet. We anticipate that the development of stronger open-source models will reduce researchers’ dependence on paid services of proprietary models.
@@ Appendix A Additional Experiments
A.1 Data Annotation with Open-Source Models
The RefAug data used in the main experiments are annotated by GPT-4-turbo. However, such annotation requires paid API calls to OpenAI’s service, which comes with a restrictive license. To this end, we explore whether state-of-the-art open-source models can also serve as data annotators. We employ the recently released LLaMA-3-70B-Instruct model Meta (2024) for data annotation using the same prompt shown in Figure 8, and train a Mistral-7B model based on this data. According to results in Table 9, RefAug data annotated by LLaMA-3-70B-Instruct yields a similar improvement in Mistral’s performance on standard math reasoning tasks. However, the reflective reasoning capability of the resulting model falls short of its counterpart trained with GPT-annotated data. This suggests that developing models with advanced reflective math reasoning skills demands higher quality data than what is typically required for standard forward reasoning in single-round QA.
A.2 Results on LLaMA-3
@@
Data GSM MATH Mathematics MAWPS SVAMP MMLU-Math SAT-Math Avg.
Original Solutions 56.25 13.96 14.80 73.07 53.50 37.68 31.82 40.15
GPT-4-turbo Solutions 65.73 23.10 23.90 81.14 68.80 40.25 41.36 49.18
+ RefAug 71.80 26.12 29.50 82.84 70.80 44.76 57.73 54.79
Table 12: Comparison between using synthetic solutions written by GPT-4-turbo and using the originally annotated ones in GSM8k and MATH training sets, as well as applying RefAug on the synthetic solutions. Solutions written by GPT-4-turbo are of much higher quality than the original ones.
A.3 Gemma on Reflective Math Reasoning
A.4 Quality of GPT-Written Answers
@@
Dataset Source Target Overlap
GSM8k Train Question Test Question 1
Train Answer Test Answer 0
RefAug Test Answer 0
MATH Train Question Test Question 228
Train Answer Test Answer 167
RefAug Test Answer 5*
A.5 Risk of Data Contamination
A.6 Training and Inference Efficiency
@@
Appendix B Detailed Task Settings
B.1 Standard Math Reasoning
Datasets
Training Settings
@@ Appendix E RefAug Annotation Prompt
Appendix F License of Artifacts
All training and evaluation datasets used in our experiments are publicly accessible and have been used in many prior researches. We note that the collection of RefAug data, if annotated by an external model, should comply with its terms of use. For example, using GPT-generated data is subject to the terms of use of OpenAI services [^5] https://openai.com/policies/terms-of-use/, and using LLaMA-generated data is subject to Meta’s LLaMA license agreement [^6] https://llama.meta.com/llama3/license/.

### v2

@@ Abstract
Zhihan Zhang1thanks:   This work was done when Zhihan, Zhenwen, and Mengzhao were interns at Tencent AI Lab, Seattle., Tao Ge2, Zhenwen Liang1†, Wenhao Yu2, Dian Yu2, Mengzhao Jia1†, Dong Yu2, Meng Jiang1 1University of Notre Dame    2Tencent AI Lab, Seattle zzhang23@nd.edu
@@ 3 Approach
Data Annotation  Following a common approach Li et al. (2023); Yu et al. (2024); Li et al. (2024a), we employ an expert LM, GPT-4-turbo, to annotate the reflective sections for high-quality reasoning paths and minimal human effort [^2] We also tried LLaMA-3-70B for data annotation in § 4.5 but its performance lags behind GPT-4-turbo.. This entails reviewing the original problem and solution to generate a section consisting of the aforementioned two types of reflective reasoning. We prompt the expert model to choose between abstraction and analogy in follow-up reasoning based on the problem context. Figure 2 shows an annotated example with alternative reasoning and follow-up analogy, and the full annotation prompt is in Appendix E.
@@ 4.2.1 Tasks
Table 3: MathChat results compared with other open-source 7B math models. Baseline scores are from Liang et al. (2024). The best scores are bolded and the second bests are underlined. GPT models are listed as a reference for state-of-the-art performance. *Including both supervised fine-tuning and reinforcement learning data.
@@ 4.4.5 Stability of RefAug Data Annotation
Data GSM MATH Mathematics MAWPS SVAMP MMLU SAT Avg. FQA-2nd FQA-3rd EC
Standard 56.25 13.96 14.80 73.07 53.50 37.68 31.82 40.15 25.72 15.25 50.68
+ RefAug (GPT) 60.05 17.36 19.40 80.25 59.30 43.63 48.64 46.95 35.36 27.54 72.99
+ RefAug (LLaMA) 62.02 17.00 17.80 80.29 61.60 39.43 44.55 46.10 32.63 23.90 50.00
Table 9: Training Mistral-7B with data where reflection sections are annotated by GPT-4-turbo or LLaMA-3-70B-Instruct. Data annotated by LLaMA-3 yields similar improvements in standard math reasoning tasks, but fails to match GPT-annotated data in enhancing Mistral’s reflective reasoning capabilities.
4.5 Data Annotation with Open-Source Models
Besides using GPT to annotate RefAug data, we explore whether state-of-the-art open-source models can also serve as data annotators. We employ LLaMA-3-70B-Instruct Meta (2024) for data annotation using the same prompt shown in Figure 8, and train a Mistral-7B model based on this data. According to results in Table 9, RefAug data annotated by LLaMA-3 yields a similar improvement in Mistral’s performance on standard math reasoning tasks. However, the reflective reasoning capability of the resulting model falls short of its counterpart trained with GPT-annotated data. This suggests that developing models with advanced reflective math reasoning skills demands higher quality data, compared to what is typically required for standard forward reasoning in single-round QA.
4.6 Data Contamination Analysis
In addition to the above perspectives, further analyses of RefAug’s impact on model efficiency are presented in Appendix A.5.
Dataset Source Target Overlap
GSM8k Train Question Test Question 1
Train Answer Test Answer 0
RefAug Test Answer 0
MATH Train Question Test Question 228
Train Answer Test Answer 167
RefAug Test Answer 5*
@@ 5 Conclusion
Some previous data augmentation studies in math reasoning created millions of data instances with OpenAI’s GPT models Li et al. (2024a); Tang et al. (2024); Huang et al. (2024b). While testing our method at a similar scale would be valuable, budget constraints limit our ability to do so. For instance, our augmentation data for MetaMath is capped at 40K instances. In §4.5, we note that LLaMA-3-70B shows some promising performance in annotating RefAug data for math reasoning tasks, though its capabilities have not fully matched those of GPT-4 yet. We anticipate that the development of stronger open-source models will reduce researchers’ dependence on paid services of proprietary models.
Acknowledgements
We would like to thank Hongming Zhang (Tencent AI Lab) for his valuable suggestions on experimental design and paper writing. We also thank Fangkai Jiao (Nanyang Technological University) and Zhenyu Wu (Xi’an Jiaotong University) for their suggestions that help shape our idea.
@@ Appendix A Additional Experiments
A.1 Results on LLaMA-3
@@ A.2 Gemma on Reflective Math Reasoning
A.3 Quality of GPT-Written Answers
@@
Dataset Alternative Follow-up
GSM8K 96% 96%
MATH 76% 72%
Table 14: The percentage of error-free RefAug annotations by GPT-4-turbo, including the alternative reasoning section and the follow-up reasoning section.
@@
A.4 Quality of GPT-annotated Reflective Sections
We analyze the correctness of GPT-annotated reflective sections by manually reviewing 50 samples (25 from GSM8K, 25 from MATH) in the training set. The results, as shown in Table 14, indicate that generating reflective sections is generally easier for GPT than solving entirely new problems. This is due to the fact that we provide both the original problem and solution during RefAug annotation. Consequently, the correctness of the annotated reflective sections is generally satisfactory.
Verification of LM-generated data is a common challenge in data augmentation. We did not dive deep into answer verification in this paper for two reasons: (1) Common methods like self-consistency voting or LM-based validation are orthogonal to our study’s focus on different augmentation types. (2) (2) Studies have indicated that data verification often does not lead to significant performance gains, and noisy answers could help training as well Yu et al. (2024); Tang et al. (2024); Li et al. (2024a). This is because such answers often include many correct reasoning steps before making an error, and filtering them trades data diversity for correctness.
A.5 Training and Inference Efficiency
@@
Appendix B Detailed Task Settings
B.1 Standard Math Reasoning
Datasets
Training Settings
@@ Appendix E RefAug Annotation Prompt
Appendix F License of Artifacts
We note that the collection of RefAug data, if annotated by an external model, should comply with its terms of use. For example, using GPT-generated data is subject to the terms of use of OpenAI services [^5] https://openai.com/policies/terms-of-use/, and using LLaMA-generated data is subject to Meta’s LLaMA license agreement [^6] https://llama.meta.com/llama3/license/.